1. a) Two IPod salesmen A and B must each make two calls per day, one in the morning and in the afternoon. A has a probability 0.4 of selling a laptop on call, while B has a probability 0.1 of a sale. A works independently of B, and for each salesman, morning and afternoon are independent of each other.

Find the probability that, in one day:


i) A sells two iPods [2 Marks]


ii) A sells just one iPod [2 Marks]


iii) B makes at least one sale [3 Marks]


iv) Between them A and B make exactly one sale. [4Marks]




1
Expert's answer
2021-12-13T11:10:38-0500

i)


"P(AA)=0.4(0.4)=0.16"


ii)


"P(A\ only\ 1)=P(AA^C)+P(A^CA)"

"=0.4(1-0.4)+(1-0.4)(0.4)=0.48"

iii)


"P(B\ at\ least\ 1)=1-P(B^CB^C)"

"=1-(1-0.1)(1-0.1)=0.19"

iv)


"P(only\ 1)=P(A\ only\ 1)P(B^CB^C)"

"+P(B\ only\ 1)P(A^CA^C)"

"=(P(AA^C)+P(A^CA))P(B^CB^C)"

"+(P(BB^C)+P(B^CB))P(A^CA^C)"

"=(0.4(1-0.4)+(1-0.4)(0.4))(1-0.1)(1-0.1)"

"+(0.1(1-0.1)+(1-0.1)(0.1))(1-0.4)(1-0.4)"

"=0.4536"